Calculate P(X < 8) if:
(a) X~Unif[5;10] (b) X~Exp(0:5)
(c) X~N(10;4) (d) X~LN(2;4)
Answer #1

a)

X~Uni(5,10). So, f(x)=rac{1}{b-a}=rac{1}{10-5}=0.2,5leq xleq 10

P(X<8)=int_{5}^{8}f(x)dx=int_{5}^{8}0.2dx=0.2[x]_{5}^{8}=0.2[8-5]=0.6

b)

X~exp(0.5). So, -0.5

P(X < 8) = F(8) = 1-e-0.5(8) = 0.9817

c)

X~N(10,4)

8-10 PlX < 8) = P(Z <--) = P(Z <-1) = P(Z > 1) = 1-P(Z < 1) 1-0.8413= 0.1587

d)

X~LN(2,4)

So, Ln(X)~N(2,4)

P(X<8)=P(Ln(X)<Ln(8))=P(Z<rac{Ln(8)-2}{sqrt{4}})=P(Z<0.04)=0.5160